jar contains 10 red marbles numbered 1 to 10 and 8 blue marbles numbered 1 to 8. A marble is drawn at random from the jar. Find the probability of the given event.

(a) The marble is red
Your answer is :

(b) The marble is odd-numbered
Your answer is :

(c) The marble is red or odd-numbered
Your answer is :

(d) The marble is blue or even-numbered
Your answer is :
1
Expert's answer
2020-06-30T14:54:11-0400

 the probability of the given event.

(a) The marble is red

answer is : 10/18=5/9

(b) The marble is odd-numbered

answer is : (5+4)/18=9/18=0.5

(c) The marble is red or odd-numbered

answer is : (10+9-5)/18=14/18=7/9

(d) The marble is blue or even-numbered

answer is : (8+9-4)/18=13/18